6. Carla invested $14,000, part at 9% and part at 12%. If the total interest at the end of the year is $1,290, how much did she invest at 9%?
A) $1,000
B) $14,000
C) $13,000
D) $12,000
no book Thanks so very very much fro your time!!

Answer by checkley71(8403)

.09X+.12(14000-X)=1290
.09X+1680-.12X=1290
-.03X=1290-1680
-.03X=-390
X=-390/-.03
X=13000 AMOUNT INVESTED @ 8%
PROOF
.09*13000+.12(14000-13000)=1290
1170+.12*1000=1290
1170+120=1290
1290=1290